OpenStudy (anonymous):

Function f is one-to-one. find it's inverse and check your answer. I did that and got to 4y=2-x, then I don't know what to do next

ganeshie8 (ganeshie8):

whats your original function ?

OpenStudy (anonymous):

f(X)=4x+2

ganeshie8 (ganeshie8):

replace f(x) by y : \[y = 4x+2\]

OpenStudy (anonymous):

Then switch the x and y

ganeshie8 (ganeshie8):

yes \[x = 4y+2\]

ganeshie8 (ganeshie8):

then solve y !

ganeshie8 (ganeshie8):

thats it

OpenStudy (anonymous):

Do I divide the four by the 4y, 2, and x?

ganeshie8 (ganeshie8):

\[x = 4y+2\] you need to isolate \(y\), first you subtract \(2\) both sides \[x-2 = 4y\]

ganeshie8 (ganeshie8):

then dividing 4 both sides isolates the y completely : \[\dfrac{x-2}{4} = y\]

ganeshie8 (ganeshie8):

which is same as writing \[y = \dfrac{x-2}{4} \]

ganeshie8 (ganeshie8):

thats the inverse function ^

OpenStudy (anonymous):

Thank you